OpenStudy (anonymous):

i will give u a mettle if u answer the question 9. Peter is trying to buy fencing for the perimeter of his garden. His garden is in the shape of a rectangle with a length of 2(x+6)feet and a width of 3.5x feet. How many feet of fencing will he need to buy? Write and simplify an expression to represent this situation? What properties did you use?

OpenStudy (mathstudent55):

What is the perimeter of a rectangle?

OpenStudy (anonymous):

P=2(l+w)

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

A rectangle has 4 sides: 2 lengths and 2 widths. What is the perimeter of a rectangle? How do you calculate it if you know the lengths of the 4 sides?

OpenStudy (anonymous):

l=14cm and w= 8cm

OpenStudy (mathstudent55):

You have the correct formula above for the perimeter: P = 2(L + W) Here, the length, L = 2(x + 6), and the width, W = 3.5x Plug those values into the formula to get: P = 2[2(x + 6) + 3.5x] Now simplify the expression.